OpenStudy (annie96):

Which ordered pair is in Quadrant III? (Points : 4) (–6, –4) (–2, 7) (0, 5) (1, 9) What is the slope of a line perpendicular to the line with equation y = 4x + 5? (Points : 4) 4 –4 ( -1/4) –5 What is the slope of a line parallel to the line with equation 5x – 2y = 11? (Points : 4) (-11/2) (-5/2) (-2/5) (5/2)

OpenStudy (annie96):

Which of the following has a slope of –2 and a y-intercept of 4? (Points : 4) y = –2x + 4 y = 2x + 4 y = 2x – 4 y = –2x – 4 Question 6.6. Which ordered pair is on the graph of the equation 3x + 7y = 11? (Points : 4) (1, –2) (0,–4) (6, –1) (0, 4)